OpenStudy (anonymous):

Find derivative of cos 2x using chain rule? Please help, studying for midterm?

OpenStudy (diamondboy):

do u understand d chain rule?

OpenStudy (anonymous):

I understand the chain rule. Would you get (-sin x) (2x) (2)?

OpenStudy (diamondboy):

yep

OpenStudy (diamondboy):

wait its -sin2x

OpenStudy (anonymous):

Okay the problem I'm having is with simplification.. what do I do from here?>

OpenStudy (diamondboy):

just put d 2 in front of sin2x

OpenStudy (anonymous):

I thought chain rule was f' x g x g'

OpenStudy (diamondboy):

so its -2sin2x

OpenStudy (anonymous):

(-sin x) (2x) (2) simplifies to -2sin2x?

OpenStudy (solomonzelman):

\[\frac{d}{dx} \left( \cos(2x) \right) = -\sin(2x) \times \frac{d}{dx} \left(2x \right)\]

OpenStudy (solomonzelman):

yes.

OpenStudy (diamondboy):

yep

OpenStudy (anonymous):

(-sin x) (2x).. why wouldn't this be sin (2x^2)?

OpenStudy (diamondboy):

cause (-sin2x)(2) u cannot separate the 2x from sin they are married for life